OpenStudy (anonymous):

Sketch the region enclosed by the given curves. Decide whether to integrate with respect to x or y. Draw a typical approximating rectangle and label its height and width. 6x+y^2=13, x=2y find area of the region.

OpenStudy (amistre64):

x = -(1/6)y^2 + 13/6 ; x = 2y

OpenStudy (amistre64):

id integrate with respect to y in the end; maybe :)

OpenStudy (anonymous):

when you integrate with respect to y, do you have the it in terms of x or y ? im a little confused D:

OpenStudy (amistre64):

terms of y

OpenStudy (amistre64):

we need to equate the 2 equations i wrote and find the 'boundarys'

OpenStudy (amistre64):

(-1/6)y^2 -2y +13/6 = 0 will give us the interval between the ys

OpenStudy (amistre64):

multiply by -6 to normalize this thing right? y^2 +12y -13 = 0

OpenStudy (anonymous):

i think it should be integrated wrt x.check the curve please...which i have given...

OpenStudy (amistre64):

y = -13, and y = 1

OpenStudy (amistre64):

coulda tried to save it as a jpeg ya know :) loads faster

OpenStudy (amistre64):

wrt x is fine; but trickier; wrt y is easier and more straight forward

OpenStudy (anonymous):

for the region left where x=2y intersects the parabola. the y is x/2, and for the right the y=sqrt(13-6x).

OpenStudy (amistre64):

when we integrate this area, we need to subtract one function from the other to get the total area; do we do absolute values? or let the areas lie where they are either pos or neg?

OpenStudy (anonymous):

so the integral would look something like this.. \[\int\limits_{-13}^{1} (-1/6)y^2 +13/6-2y)dy\]

OpenStudy (amistre64):

i believe so :)

OpenStudy (anonymous):

than you both so much :)

OpenStudy (anonymous):

thank*

OpenStudy (amistre64):

if it turns out negative; just take the absolute value; it simply means you subtracted in the wrong order for example: 5-3 = 2 3-5 = -2; but |-2| = 2

OpenStudy (anonymous):

gotcha. thanks a bunch :)

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

it all worked out. :D you rock :)

OpenStudy (amistre64):

:) thnx